OpenStudy (anonymous):

use the subtraction to solve system of equation x+2y=12 -x=-y-6 a) (9,3) b) ( 6,0) c ( 8,2) d) ( 7,1)

OpenStudy (anonymous):

x+2y=12 -x=-y-6 The easiest way of solving this is to add the two equations together. This will result x to drop. x+2y-x=12-y-6 3y=6 y=2 plug into one of the equations x+2*2=12 x=8 Solution: (8.2)

OpenStudy (anonymous):

thanks